OpenStudy (k_lynn):

The perimeter of a rectangle is 68 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 9 : 8. Which of the following would be the best equation to use to solve this problem?

OpenStudy (k_lynn):

Here it is.

OpenStudy (k_lynn):

@iambatman

OpenStudy (k_lynn):

can you help me?

OpenStudy (k_lynn):

@e.mccormick would you be able to help me with this?

OpenStudy (anonymous):

What is the perimeter of a rectangle?

OpenStudy (k_lynn):

68ft.

OpenStudy (anonymous):

what is the equation look like

OpenStudy (k_lynn):

Well for this one at least.

OpenStudy (anonymous):

Do you know what a perimeter is

OpenStudy (k_lynn):

ya, the amount of all the sides of a shape. Say all the sides of a square were 4in. the perimeter would be 16in.

OpenStudy (anonymous):

Ok so make an equation for a rectangle,

OpenStudy (k_lynn):

I'm just not sure how to use an equation to solve it.

OpenStudy (k_lynn):

I'm thinking its D. But I'm not entirely sure.

OpenStudy (anonymous):

Hint:  what is the perimeter of this?

OpenStudy (k_lynn):

l*2+w*2

OpenStudy (k_lynn):

Well now, I'm kind of thinking its B.

OpenStudy (anonymous):

It's actually just 2l+2w = 68 haha

OpenStudy (k_lynn):

Ok, so A then?

OpenStudy (anonymous):

Bingo

OpenStudy (k_lynn):

Ok, thank you.

OpenStudy (anonymous):

Don't let the ratio confuse you ;)

OpenStudy (k_lynn):

I'll try not to.

OpenStudy (e.mccormick):

Actually, the ratio is not there to confuse you. You are trying to find the actual dimensions, which the ratio lets you do.

OpenStudy (e.mccormick):

A only works if it is a square. It is not a square, therefore A is wrong. B is a fraction so would have nothing to do with a paremeter, so wrong. C does not account for there being 2 of each side, so wrong. D, on the other hand, does account for the ratio and four sides.

OpenStudy (k_lynn):

Oh, ok thanks. That really clears some things up.

OpenStudy (e.mccormick):

9x+9x+8x+8x=68 18x+16x=68 18x+16x=68 34x=68 x=2 Therefore, the long side is 18 and the short side is 16. I can confirm by going back to the base formula: 2l+2w=68 2(18)+2(16)=68 36+32=68 68=68 None of the others are actually able to do that.

OpenStudy (anonymous):

Well that is true you can do 9x+9x+8x+8x=68 =>34x=68 => x = 2 and use this to check the for parameter haha.

OpenStudy (k_lynn):

Thank you, that's really helpful. I think I get this now.

OpenStudy (e.mccormick):

Yah, the ratio was there as information. However, the fraction was there to confuse you because they used the word ratio and the two are somewhat interchangable.

OpenStudy (e.mccormick):

A looks a lot like the actual formula, 2l+2w=P, but note they used x only and not x and y, so it is really the perimeter of a square! 4x=P for a square. C was similarly flawerd because it had the ratio but not the 2!